OpenStudy (anonymous):

Find antiderivative 5x-x^(1/2) How come on the -x side the answer is 2x^(3/2) / 3 ? Does the 3/2 get flipped when it goes to the bottom?

OpenStudy (jamesj):

Because \[ \frac{1}{3/2} = \frac{2}{3} \]

OpenStudy (anonymous):

I wish I knew its bad retricethough, I think its a monkey eagle or vietnam eagle

OpenStudy (anonymous):

Ok james so I don't just take the exponent and add one to it then put it on the bottom I actually have to take the exponent over 1?

OpenStudy (jamesj):

Remember that the antiderivative of \( x^n, n \neq 1 \) is \[ \frac{1}{n+1}x^{n+1} \ \ (+ C) \] In this case,y ou want the antiderivative of \( x^{1/2} \), hence n = 1/2. Using the rule just written down, we have n = 1/2 and thus the antiderivative is \[ \frac{1}{1+ 1/2} x^{1 + 1/2} = \frac{1}{3/2} x^{3/2} = \frac{2}{3} x^{3/2} \]

OpenStudy (anonymous):

if I take x^2 for example, its an easy x^(3) / 3

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

thanks JamesJ!